hartnn (hartnn):

Check my answer! integral from 0 to infinity e^(-2t) sinht sin t/ t dt I got 1/2 (π/4-cot^(-1)3 )

OpenStudy (anonymous):

mind adding in parentheses so we can clarify your notation?

hartnn (hartnn):

\(\Large \int \limits_0^\infty e^{-2t} \sinh t \dfrac{\sin t}{ t} dt = \dfrac{1}{2}(\dfrac{\pi }{4}-\cot^{-1}3)\)

hartnn (hartnn):

need more clarification ?

OpenStudy (anonymous):

no, looks good. one moment

hartnn (hartnn):

used laplace transform L[sin t/t] = cot^-1 (s)

OpenStudy (anonymous):

ya that's what i got with that transform. gj

hartnn (hartnn):

its exactly correct?